# The First Search for Optical Transient as a Counterpart of a Month-timescale IceCube Neutrino Multiplet Event

### v1

@@ 3.5 Multi-epoch analysis
The transient source search for ICT-MJD59015 is facilitated by the ZTF observations at multiple epochs. When any single observation meets the detection-wise criteria, level 1 and 2 criteria are fulfilled. The incorporation of additional epochs into the counterpart search process enhances the resultant TPR, but it also raises the probability of contamination of background events. It is imperative to judiciously optimize the number of epochs contributing to the search to ensure the efficacy of the methodology.
We trace the epoch dependence of TPR and the number of background events by applying the level 1 and level 2 selections to the ME1 control data sample, which contains 11 observing epochs. Each of the observation epochs has a different depth due to the night sky conditions as shown in Figure 5, and we search for the best combination of epochs participating in the source search. The specific 6 epochs have been chosen to give an excellent TPR ($0.983$ in typical magnitude range, 18.5-19.0 mag) while suppressing the number of penetrating background events ($\lesssim 0.1~{}{\rm deg^{2}}$). We then apply the selections to the ME2 control data sample, which contains exactly the 6 selected epochs, but data from larger areas of the sky to gain statistical power.
The cumulative number of background objects is 67 objects, which corresponds to a background rate of $0.084\pm 0.010~{}{\rm objects~{}deg^{-2}}$. The IceCube localization area ($3\sigma$) covered by the ZTF observation is 0.41 $\rm deg^{2}$ (Figure 1). Therefore, the background rate in the final analysis for the localization area of ICT-MJD59015 is $0.034\pm 0.004$.
@@
The dominance of SNe as background objects means that the background rate may vary depending on the direction, reflecting the variation in the number of SNe due to the inhomogeneous distribution of galaxies at low redshift. This effect may be estimated by integrating the luminosity of the galaxies toward each direction. The SN rate can be approximated by the total star formation rate, and the star formation rate is roughly proportional to the $B$-band luminosity (R. C. Kennicutt, 1998). By using the GLADE+ catalog (G. Dálya et al., 2022), we estimate the integrated $B$-band luminosity ($L_{\rm B}$) of galaxies per unit area both for the region used for background analysis and the localization area of ICT-MJD59015. We find that the $L_{\rm B}$ in the localization area is 1.26 times higher than the median of the background region. The standard deviation of $L_{\rm B}$ in the background regions is 0.089 dex. For a conservative estimation of the background, we define a correction factor of 1.55, which is the ratio between $L_{\rm B}$ in the localization area and $L_{\rm B}$ at $1\sigma$ below the median value of the background regions. We scale the background rate by this factor, resulting in the final background rate of $0.053\pm 0.006$, which is used in the signal model test  [^5] Most of background SNe are located at $z=0.05-0.1$ while GLADE+ catalog is highly incomplete at such a redshift range. Thus, the derived variation of $L_{\rm B}$ in the background region is overestimated due to a high fraction of nearby galaxies. Thus, the correction factor by using $L_{\rm B}$ at 1 $\sigma$ below the median value gives a very conservative estimate of the background rate. .
@@ 3.6 Signal model tests
We discuss our statistical analysis method to test our signal model. We look for transient objects in the neutrino localization area, and the observable is the number of the transients that pass through our event selections, $n_{T}$ (see Section 3.4). Our signal hypothesis is that the neutrino-emitting object has a set of $(L_{\rm pk},~{}t_{\rm decay},~{}\Delta t)$. We estimate the probability that ZTF detects the object to be
[…]
where $f_{\rm cover}$ is the coverage of the neutrino error region and $p_{\rm fail}(z)=\prod_{i=1}^{6}(1-{\rm TPR_{\rm 1}({\it t_{i},z})TPR_{2}}(t_{i},z))$ is the probability that ZTF fails to detect the object at a distance $z$ at an epoch $t_{i}$. ${\rm TPR_{obj}=TPR_{3}TPR_{4}}$ is the TPR of object-wise selections (levels 3 and 4). Here ${\rm TPR}_{n}$ denotes the TPR of $n$th selection process. The sensitivity of the ZTF data depends on the epoch $t_{i}$, and we set ${\rm TPR_{1}}=0$ when the flux by our signal model is lower than the sky noise level [^6] Since we choose the 6 epochs that have deep enough limiting magnitudes, this treatment does not affect our resulting constraint maps.. In addition, we count the number of epochs that satisfy the condition that the signal flux is larger than the sky noise level, and set $p_{\rm fail}=1$ if the number count is less than or equal to 1. We should note that $10<n_{\rm sig}<$ 1 is always satisfied. Given the background event, $\mu_{B}$, which should follow the Poisson distribution, we define likelihood as
=$\mathcal{L}=\left\{\begin{array}[]{ll}e^{-\mu_{B}}(1-n_{\rm sig}){}{}&~{}~{}(n_{T}=0)\\ 1-e^{-\mu_{B}}(1-n_{\rm sig}){}{}&~{}~{}(n_{T}\geq 1)\end{array}\right.,$ (6)
@@
where $(\hat{L}_{\rm pk},~{}\hat{t}_{\rm decay}~{},\hat{\Delta t})$ are chosen so that $\mathcal{L}$ is maximized.
@@ 5.1 Implications for TDE models
TDEs have been discussed as high-energy neutrino sources since the discovery of the jetted TDEs in 2011 (X.-Y. Wang et al., 2011; C. Lunardini & W. Winter, 2017; L. Dai & K. Fang, 2017). However, it turned out that they are too rare to be the dominant source of cosmic neutrino background observed on Earth; Non-detection of IceCube lower-energy neutrinos constrain the rare population of neutrino sources with $R_{0}\lesssim 10^{-7}\rm~{}Mpc^{-3}~{}yr^{-1}$ (R. Abbasi et al., 2023), which strongly disfavors the jetted TDEs ($R_{0}\lesssim 3\times 10^{-10}\rm~{}Mpc^{-3}~{}yr^{-1}$; G. C. Brown et al. 2015).
Our analysis disfavors long-duration transients as neutrino sources, especially for $t_{\rm decay}\gtrsim 50$ days. However, this does not mean that we could disfavor TDEs as the dominant source of the cosmic neutrino background. All the candidate TDE-neutrino associations reported so far are non-jetted TDEs and have long time lags between the optical peak and neutrino detection, typically $\gtrsim$ 100 days (R. Stein et al., 2021; S. Reusch et al., 2022; N. Jiang et al., 2023; S. van Velzen et al., 2024; C. Yuan et al., 2024). This feature could be explained by some of current scenarios, such as collisionless corona formation (K. Murase et al., 2020), wind-cloud interactions (H.-J. Wu et al., 2022), and choked delayed jets (J.-H. Zheng et al., 2023; M. Mukhopadhyay et al., 2024a). If such a delayed neutrino emission is typical, a TDE is unlikely to be the origin of triplet events with a 30-day time window, as the expected duration of the neutrino emission is much longer than the time window of the multiplet alert. In this case, we cannot constrain the TDE-neutrino paradigm by this analysis.
@@ 5.2 Implications for SLSN models
Although SLSNe have not been discussed as high-energy neutrino sources frequently, they are good candidates of high-energy neutrino sources (see T. Pitik et al. 2022 for possible association between IC-200530 & AT2019fdr). The power source of SLSNe are still uncertain (T. J. Moriya, 2024), and three models are actively discussed: radioactive decay of a large amount of 56Ni, CSM interaction, and energy injection by a central remnant. Although high-energy neutrinos are unlikely to be produced for the radioactive-decay scenario, the other two power-sources have been actively discussed as high-energy neutrino sources in the context of core-collapse SNe (K. Murase et al., 2011; V. N. Zirakashvili & V. S. Ptuskin, 2016; M. Petropoulou et al., 2017; K. Fang et al., 2020; S. S. Kimura & T. J. Moriya, 2024) and neutron-star mergers (K. Fang & B. D. Metzger, 2017; M. Mukhopadhyay et al., 2024b). Both of these scenarios predict that the time window for neutrino emission is about $10-30$ days, which is a good match with the multiplet alert.
The main differences of neutrino signals in these two models are their expected neutrino spectra: CSM interaction SNe produce neutrinos via $pp$ channel with shock-accelerated CRs, which leads to the canonical spectra of $dN/dE_{\nu}\sim E_{\nu}^{-2}$ or softer. On the other hand, magnetar engines produce neutrinos via the $p\gamma$ channel with CRs accelerated at polar-cap or reconnection, which make the neutrino spectra much harder than the canonical one, typically $dN/dE_{\nu}\sim E_{\nu}^{-1}$ with the spectral peak at EeV energies. Considering these features, central-engine SLSNe are likely to produce multiplet signals with PeV-EeV energies. Thus, the current multiplet alert with TeV-PeV energies should be most likely associated with the interaction-powered SLSNe.
Our null result disfavors SLSNe as the dominant source of cosmic neutrino background. This is consistent with the theoretical expectation based on an energetics argument shown below. CSM interaction features are observed not only SLSNe but also SNe IIn that are roughly 100 times more frequent but 10 times less luminous. This implies that SN IIn have 10 times higher energy budget than SLSNe, and thus, the cosmic neutrino background should be dominated by SN IIn as long as both SN IIn and SLSNe share the same energy release mechanism through the CSM interaction. Although current neutrino detectors are not sensitive enough to detect multiplet signals from SN IIn, multiplet alerts with future detectors would be able to constrain this type of transients as cosmic neutrino sources.
@@ 6 Summary
Appendix A Sky regions for Background analysis
For the background modeling (Section 3.3), we use control data samples from three fields: the Wide Field (WF) sample, the Multiple Epoch 1 (ME1) sample, and the Multiple Epoch 2 (ME2) sample. The WF sample consists of all alerts in the $g$-band on July 30, 2020, excluding the region around the direction of the IceCube triplet event ICT-MJD59015. Figure 8 shows the sky distribution of the alerts of the WF sample passing each selection level. For the multiple-epoch analysis, we use data from sky regions that were repeatedly observed over 11 epochs (from July 14 to 31, 2020) for the ME1 sample and 6 epochs (from July 14 to 23, 2020) for the ME2 sample (Figure 9).
@@ Appendix A Sky regions for Background analysis
Figure 9: The sky region used in our multi-epoch analysis. Red regions represent the area repeatedly observed from July 6 to 31, 2020. For 6 epochs analysis, we add blue regions which are repeatedly observed from July 14 to July 23. The gray region represents the field which includes IceCube localization area of the triplet ICT-MJD59015.

### v2

@@ 3.5 Multi-epoch analysis
The transient source search for ICT-MJD59015 is facilitated by the ZTF observations at multiple epochs. When any single observation meets the detection-wise criteria, level 1 and 2 criteria are fulfilled. The incorporation of additional epochs into the counterpart search process enhances the resultant TPR, but it also raises the probability of contamination of background events (see also Appendix B and Figure 10). It is imperative to judiciously optimize the number of epochs contributing to the search to ensure the efficacy of the methodology.
We trace the epoch dependence of TPR and the number of background events by applying the level 1 and level 2 selections to the ME1 control data sample, which contains 11 observing epochs. As we mentioned in Section 2.2, we do not include the shallow observations on July 6, 2020, since the data on that day are not effective to detect dimmer objects and just increase the number of background objects. Each of the observation epochs has a different depth due to the night sky conditions as shown in Figure 5, and we search for the best combination of epochs participating in the source search. The specific 6 epochs have been chosen to give an excellent TPR ($0.983$ in typical magnitude range, 18.5–19.0 mag) while suppressing the number of penetrating background events ($\lesssim 0.1~{\rm deg^{2}}$). We then apply the selections to the ME2 control data sample, which contains exactly the 6 selected epochs, but data from larger areas of the sky to gain statistical power.
The cumulative number of background objects is 67 objects, which corresponds to a background rate of $0.084\pm 0.010~{\rm objects~deg^{-2}}$. For comparison, applying the ZTF standard filtering results in a background rate of $0.175\pm 0.015~{\rm objects~deg^{-2}}$. The IceCube localization area ($3\sigma$) covered by the ZTF observation is 0.41 $\rm deg^{2}$ (Figure 1). Therefore, the background rate in the final analysis for the localization area of ICT-MJD59015 is $0.034\pm 0.004$.
@@
The dominance of SNe as background objects means that the background rate may vary depending on the direction, reflecting the variation in the number of SNe due to the inhomogeneous distribution of galaxies at low redshift. This effect may be estimated by integrating the luminosity of the galaxies toward each direction. The SN rate can be approximated by the total star formation rate, and the star formation rate is roughly proportional to the $B$-band luminosity (R. C. Kennicutt, 1998). By using the GLADE+ catalog (G. Dálya et al., 2022), we estimate the integrated $B$-band luminosity ($L_{\rm B}$) of galaxies per unit area both for the region used for background analysis and the localization area of ICT-MJD59015. We find that the $L_{\rm B}$ in the localization area is 1.26 times higher than the median of the background region. The standard deviation of $L_{\rm B}$ in the background regions is 0.089 dex. For a conservative estimation of the background, we define a correction factor of 1.55, which is the ratio between $L_{\rm B}$ in the localization area and $L_{\rm B}$ at $1\sigma$ below the median value of the background regions. We scale the background rate by this factor, resulting in the final background rate of $0.053\pm 0.006$, which is used in the signal model test  [^5] Most of background SNe are located at z=0.050.1z=0.05-0.1 while GLADE+ catalog is highly incomplete at such a redshift range. Thus, the derived variation of $L_{\rm B}$ in the background region is overestimated due to a high fraction of nearby galaxies. Thus, the correction factor by using $L_{\rm B}$ at 1 $\sigma$ below the median value gives a very conservative estimate of the background rate. .
@@ 3.6 Signal model tests
We discuss our statistical analysis method to test our signal model. Here, we assume that the triplet is of astrophysical origin and set the signalness to unity. We look for transient objects in the neutrino localization area, and the observable is the number of the transients that pass through our event selections, $n_{T}$ (see Section 3.4). Our signal hypothesis is that the neutrino-emitting object has a set of $(L_{\rm pk},~t_{\rm decay},~\Delta t)$.
We estimate the probability that ZTF fails to detect the object at a distance $z$ at epochs $t_{i}$ as
$p_{\rm fail}(z)=\prod_{i=1}^{6}(1-{\rm TPR_{\rm 1}({\it t_{i},z})TPR_{2}}(t_{i},z)),$ (3)
where ${\rm TPR}_{n}$ denotes the TPR of $n$th selection process. The sensitivity of the ZTF data depends on the epoch $t_{i}$. Thus, when the flux by our signal model is lower than the sensitivity limit, we set ${\rm TPR_{1}}(t_{i},z)=0$  [^6] Since we choose the 6 epochs that have deep enough limiting magnitudes, this treatment does not affect our resulting constraint maps.. In addition, we count the number of epochs that satisfy the condition that the signal flux is larger than the sky noise level, and set $p_{\rm fail}=1$ if the number count is less than or equal to 1. Since $\rm TPR_{1}$ depends on the optical flux at the epoch $t_{i}$, $p_{\rm fail}$ depends on the source parameters, i.e., a higher $L_{\rm pk}$, a longer $t_{\rm delay}$, and a shorter $\Delta t$ lead to a lower $p_{\rm fail}$. Taking into accout the source redshift distribution, we estimate the probability that ZTF detects the object to be
$n_{\rm sig}={\rm TPR}_{\rm obj}f_{\rm cover}\int dz(1-p_{\rm fail}(z))P_{\rm dist}(z),$ (4)
where $f_{\rm cover}$ is the coverage of the neutrino error region and $\rm TPR_{obj}=TPR_{3}TPR_{4}$} is the TPR of object-wise selections (levels 3 and 4). $P_{\rm dist}(z)$ is the probability distribution functions of the sources of triplet events introduced in Section 3.2. We should note that $0<n_{\rm sig}<1$ is always satisfied. Given the background event, $\mu_{B}$, which should follow the Poisson distribution, we define likelihood as
$\mathcal{L}=\left\{\begin{array}[]{ll}e^{-\mu_{B}}(1-n_{\rm sig})&~~(n_{T}=0)\\ 1-e^{-\mu_{B}}(1-n_{\rm sig})&~~(n_{T}\geq 1)\end{array}\right$., (7)
@@
[…]
@@ 5.1 Implications for TDE models
TDEs have been discussed as high-energy neutrino sources since the discovery of the jetted TDEs in 2011 (X.-Y. Wang et al., 2011; C. Lunardini & W. Winter, 2017; L. Dai & K. Fang, 2017). However, it turned out that they are too rare to be the dominant source of cosmic neutrino background observed on Earth; Non-detection of IceCube lower-energy neutrinos constrain the rare population of neutrino sources with $R_{0}\lesssim 10^{-7}\rm~Mpc^{-3}~yr^{-1}$ (R. Abbasi et al., 2023a), which strongly disfavors the jetted TDEs ($R_{0}\lesssim 3\times 10^{-10}\rm~Mpc^{-3}~yr^{-1}$; G. C. Brown et al. 2015).
Our analysis disfavors long-duration transients as neutrino sources, especially for $t_{\rm decay}\gtrsim 50$ days. However, this does not mean that we could disfavor TDEs as the dominant source of the cosmic neutrino background. All the candidate TDE-neutrino associations reported so far are non-jetted TDEs and have long time lags between the optical peak and neutrino detection, typically $\gtrsim$ 100 days (R. Stein et al., 2021; S. Reusch et al., 2022; N. Jiang et al., 2023; S. van Velzen et al., 2024; C. Yuan et al., 2024). This feature could be explained by some of current scenarios, such as collisionless corona formation (K. Murase et al., 2020), wind-cloud interactions (H.-J. Wu et al., 2022), choked delayed jets (J.-H. Zheng et al., 2023; M. Mukhopadhyay et al., 2024a), and IR dust echos (W. Winter & C. Lunardini, 2023). If such a delayed neutrino emission is typical, a TDE is unlikely to be the origin of triplet events with a 30-day time window, as the expected duration of the neutrino emission is much longer than the time window of the multiplet alert. In this case, we cannot constrain the TDE-neutrino paradigm by this analysis.
@@ 5.2 Implications for SLSN models
Although SLSNe have not been discussed as high-energy neutrino sources frequently, they are good candidates of high-energy neutrino sources (See T. Pitik et al. 2022 for possible association between IC-200530 & AT2019fdr as a SLSN. However, AT2019fdr is also considered to be a TDE; cf. S. Reusch et al. 2022.). The power source of SLSNe are still uncertain (T. J. Moriya, 2024), and three models are actively discussed: radioactive decay of a large amount of 56Ni, CSM interaction, and energy injection by a central remnant. Although high-energy neutrinos are unlikely to be produced for the radioactive-decay scenario, the other two power-sources have been actively discussed as high-energy neutrino sources in the context of core-collapse SNe (K. Murase et al., 2011; V. N. Zirakashvili & V. S. Ptuskin, 2016; M. Petropoulou et al., 2017; K. Fang et al., 2020; S. S. Kimura & T. J. Moriya, 2024) and neutron-star mergers (K. Fang & B. D. Metzger, 2017; M. Mukhopadhyay et al., 2024b). Both of these scenarios predict that the time window for neutrino emission is about 103010-30 days, which is a good match with the multiplet alert.
The main differences of neutrino signals in these two models are their expected neutrino spectra: CSM interaction SNe produce neutrinos via $pp$ channel with shock-accelerated CRs, which leads to the canonical spectra of $dN/dE_{\nu}\sim E_{\nu}^{-2}$ or softer (e.g. S. S. Kimura & T. J. Moriya, 2024). On the other hand, magnetar engines produce neutrinos via the $p\gamma$ channel with CRs accelerated at polar-cap or reconnection, which make the neutrino spectra much harder than the canonical one, typically $dN/dE_{\nu}\sim E_{\nu}^{-1}$ with the spectral peak at EeV energies (e.g. M. Mukhopadhyay et al., 2024b). Considering these features, central-engine SLSNe are likely to produce multiplet signals with PeV-EeV energies. Thus, the current multiplet alert with TeV-PeV energies should be most likely associated with the interaction-powered SLSNe.
Our null result disfavors SLSNe as the dominant source of cosmic neutrino background. This is consistent with the theoretical expectation based on an energetics argument shown below. CSM interaction features are observed not only SLSNe but also SNe IIn that are roughly 100 times more frequent but 10 times less luminous. This implies that SN IIn have 10 times higher energy budget than SLSNe, and thus, the cosmic neutrino background should be dominated by SN IIn as long as both SN IIn and SLSNe share the same energy release mechanism through the CSM interaction (cf. R. Abbasi et al. 2023b). Although current neutrino detectors are not sensitive enough to detect multiplet signals from SN IIn, multiplet alerts with future detectors would be able to constrain this type of transients as cosmic neutrino sources.
@@ 6 Summary
Appendix A Signalness of neutrino multiplets
The signalness of neutrino multiplet is defined as a fraction of multiplets originating from astrophysical transient sources to the total number of multiplets with identical test statistics, $\Lambda$ (see R. Abbasi et al. (2025) for the derivation of $\Lambda$). The distribution of $\Lambda$ from astrophysical transients, $dN^{\mathrm{astro}}/d\Lambda$, is computed assuming an emission energy of neutrinos per a source, $\varepsilon_{\nu}$ and an event rate of the transients $R_{0}$ as:
$\frac{dN^{\mathrm{astro}}}{d\Lambda}=T_{\mathrm{obs}}R_{0}\int dz\,\psi(z)\,\rho(\Lambda;\mu(\varepsilon_{\nu},z))\frac{dV}{dz},$ (A1)
where $T_{\mathrm{obs}}$ is a total livetime of a dataset, dV{dV} is a volume element, $z$ is a redshift of the source, and $\psi(z)$ is an evolution of sources with redshift which is compatible with the star formation rate R. Abbasi et al. (2025). A factor of $\rho(\Lambda;\mu(\varepsilon_{\nu}))$ represents a normalized distribution of $\Lambda$ (i.e., a probability density function) when the averaged number of neutrino detection in the multiplet time window (30 days) is $\mu$, and is evaluated by the simulation of injections of signal neutrino events. The value of $\mu$ is a function of $\varepsilon_{\nu}$ and $z$, and computed based on the neutrino flux from a source at $z$ and an effective detection area of the IceCube. The distribution of $\Lambda$ from backgrounds (atmospheric neutrinos and unrelated astrophysical neutrinos) in the livetime of $T_{\mathrm{obs}}$, $dN^{\mathrm{bg}}/d\Lambda$, is evaluated by a data-driven background control sample. The signalness at $\Lambda$ is thus given by:
$\mathcal{S}(\Lambda)=\frac{\frac{dN^{\mathrm{astro}}}{d\Lambda}}{\frac{dN^{\mathrm{astro}}}{d\Lambda}+\frac{dN^{\mathrm{bg}}}{d\Lambda}}.$ (A2)
From Figure 4 (bottom panel) of R. Abbasi et al. 2025, the largest test statistic value of triplet, which corresponds to ICT-MJD59015, indicates $\mathcal{S}(\Lambda)\sim 60\%$.
Appendix B Background analysis
For the background modeling (Section 3.3), we use control data samples from three fields: the Wide Field (WF) sample, the Multiple Epoch 1 (ME1) sample, and the Multiple Epoch 2 (ME2) sample. The WF sample consists of all alerts in the $g$-band on July 30, 2020, excluding the region around the direction of the IceCube triplet event ICT-MJD59015. Figure 8 shows the sky distribution of the alerts of the WF sample passing each selection level.
As described in Section 3.3, ICT-MJD59015 localization area was observed eleven times in July 2020, except for shallower observations on July 6. Since the level 1 and 2 selections are utilized in each of the ZTF observations at multiple epochs, the TPR improves as the number of observations increases. On the other hand, the number of background objects also increases. Therefore, the epoch selection has been optimized by examining the manner in which the incorporation of multiple observations enhances the TPR in compensation of the background accumulation, using the ME1 and ME2 control samples (see also Section 3.5). For this multiple-epoch analysis, we use data from sky regions that were repeatedly observed over 11 epochs (from July 14 to 31, 2020) for the ME1 sample and 6 epochs (from July 14 to 23, 2020) for the ME2 sample (Figure 9). Figure 10 shows the relation between the TPR and the number of background objects for different number of observations. Based on analysis using ME1 sample, we chose a set of 6 epochs from July 14 to July 23 that achieves a high TPR and a low number of background objects. Then, we analyze ME2 sample to improve the statistics in estimating the number of background objects.
@@ Appendix B Background analysis
Figure 9: The sky region used in our multi-epoch analysis. Red regions represent the area repeatedly observed from July 14 to 31, 2020. For 6 epochs analysis, we add blue regions which are repeatedly observed from July 14 to July 23. The gray region represents the field which includes IceCube localization area of the triplet ICT-MJD59015.
Figure 10: Relations between the number of background objects and the TPR along with repeating observations. Red line represents measurement using the ME1 sample (July 14 to July 31) and blue line represents the measurement using the ME2 sample (July 14 to July 23) respectively. We assume that the objects are in the magnitude range of 18.5–19.0 mag to calculate the TPR.